Derive the MU function from the following TU function: π‘‡π‘ˆ = 200𝑄 βˆ’ 25𝑄 2 + 𝑄 3 From this MU function, draw up a table up to the level of Q where MU becomes negative. i.e (Q = 1, 2, etc). Graph these figures


1
Expert's answer
2021-04-28T09:33:49-0400

Answer

"Given TU= 200Q-25Q^{2}+Q^{3}"

We have to know that;


"At Q=1,\t\tTU= 200-25+1=176"

"At Q=2,\t\tTU=200\\times2-25\\times4+2^3=308"

"At Q=3,\t\tTU=200\\times3-25\\times3+3^3=402"

"At Q=4,\t\tTU=200\\times4-25\\times4^2+4^3=464"

"At Q=20,\tTU=200\\times20-25\\times20^2+20^3=2000"



When MU becomes negative TU lowers thus decreasing the level of Q.


"So MU_{n}=TU_{n}-TU_{n-1}"

"Therefore the marginal function will be ; MU=\frac{2TU}{2Q}=200-50Q+3Q^2" "at Q=1, \t\tMU=176-0=176"

"at Q=2,\t\tMU=308-176=132"

"at Q=3, \t\tMU=402-308=94"

"at Q=4,\t\tMU=464-402=62"


So graph of MU: